Barbara had money in two accounts. She had $900 more in an account paying 8% annual interest than she had in an account paying 6% annual interest. If she earned $240 in interest for the year, how much did she have invested in each account?
Answer by mananth(16946)

6% account be $ x
8% account she has x+900
..
Total interest she earned =$240
..
0.06x+0.08(x+900)=240
0.06x +0.08x + 72=240
0.14x= 240-72
0.14x= 168
/0.14
x= 168/0.14
=$1200 amount at 6 %
1200+900 = $2100 at 8%